# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With Affine Transformation Contrastive Learning

### v3

@@ Abstract
Aerial Remote Sensing (ARS) vision tasks pose significant challenges due to the unique characteristics of their viewing angles. Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications. This paper proposes the RingMo-Aerial model, aiming to fill the gap in foundation model research in the field of ARS vision. By introducing the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism and an affine transformation-based contrastive learning pre-training method, the model’s detection capability for small targets is enhanced and optimized for the tilted viewing angles characteristic of ARS. Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and effectiveness in various ARS vision tasks. Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks. This indicates the practicality and efficacy of RingMo-Aerial in enhancing the performance of ARS vision tasks.
@@ 1 Introduction
These challenges highlight the need for specialized intelligent models tailored to ARS applications. While recent foundational models in computer vision and RS have shown potential, they are often inadequate for ARS due to its unique requirements. These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
@@
In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic ARS foundational models called ARS-Adapter. Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting. ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter[6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling. This approach optimizes the performance of the ARS foundational model, making it more adaptable and effective across various ARS vision tasks.
@@
To evaluate the effectiveness of the model, we performed six downstream tasks using 13 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes. The test results show that RingMo-Aerial achieves SOTA performance on all ARS datasets and reaches the leading level on ApRS datasets, which proves its strong generalization ability for intelligent interpretation of aerial and ApRS images. Compared with the single-task model, RingMo-Aerial has apparent advantages in handling multi-tasks in different scenarios.
@@ 2.2 Visual foundation model
Visual foundation models have gained attention as general-purpose solutions for computer vision tasks. Early models, based on convolutional neural networks (CNNs) like VGG [20] and ResNet [21], used convolutional layers for feature extraction and were typically pre-trained on classification tasks using ImageNet [22]. With the advent of Transformers in natural language processing, visual Transformer models like ViT [23], and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training. In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies. Several novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed. However, these newer models have yet to achieve widespread adoption due to technical immaturity.
@@ 2.2.2 Masked Image Modeling
2.2.3 The combination of Contrastive Learning &\&& Masked Image Modeling
@@ 3 Problem Description
Figure 5: Similar targets with significant size differences in the same picture. In the Vision Transformer model, objects of different sizes occupy different numbers of tokens. The image is from the VisDrone dataset.
The unique imaging properties of ARS photography stem from the inherent properties of aerial perspective. Specifically, this perspective manifests itself in size differences within the same image due to the tilted angle of the shot. As shown in Fig.5, similar objects at different distances from the ARS are rendered at different sizes in the image.
@@
As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\alpha$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera. At different horizontal distances, the imaging size corresponding to the same target height differs. Specifically, when the target is directly beneath the camera, its height is not captured in the image. As the horizontal distance increases, the number of pixels occupied by the target’s height initially increases and then decreases. By considering the distances from the target’s lowest and highest points to the camera, given by $x=\sqrt{H^{2}+l^{2}}$ and $y=\sqrt{(H-h)^{2}+l^{2}}$ respectively, we can derive the relationship between the viewing angle $\alpha$ and the horizontal distance $l$ between the target and the camera:
Using the cosine rule:
$x^{2}+y^{2}-2xy\cos\alpha=h^{2}$ (1)
@@
$\cos\alpha=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
By setting $a=H$ and $b=H-h$, we have the target function is:
$f(l)=\cos\alpha=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
@@
Since $a=H$ and $b=H-h$, and $a>b$, it follows that $a-b=h>0$.
@@
This completes the derivation of the maximum value $\alpha$ and the corresponding minimum value $\cos\alpha$. From Eq. (2), when $l=0$, $\alpha$ is equal to 0. From Eqs.(4) and (9), $\alpha$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\alpha$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\alpha$ is again equal to 0. Assuming that the optical axis of the camera is always at the center of the target viewpoint $\alpha$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\alpha$ of the target. This indicates that if the same target is at different distances from the lens and imaging angles, it will have different representations in the image.
@@ 4.1 Method Overview
The proposed RingMo-Aerial model serves as a foundational framework with broad applicability. Its implementation involves three main stages: model design, pre-training, and fine-tuning for downstream tasks. In Section 4.2, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase. To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown in Fig.7. In Section 4.3, we combine the commonly used masked image modeling (MIM) pre-training with additional contrastive learning (CL) methods based on affine transformations during the pre-training stage. This hybrid pre-training approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown in Fig.7. Furthermore, in Section 4.4, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase. The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks.
@@ 4.2 Model Architecture
To address the challenges above, we propose the overall architecture of RingMo-Aerial. As shown in Fig.7 Pre-training Stage 1, considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
@@
Fig.7(b) illustrates the RingMo-Aerial Block structure, including two window attention mechanisms optimized by the FE-MSA mechanism. Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block. Taking the first stage as an example, the input feature vector $z^{l}$ adopts the same slicing parameters as the standard Swin-Transformer, resulting in four-channel slices $C=\{c_{1},c_{2},c_{3},c_{4}\}$. After linear normalization, it enters the FE-MSA module. For each patch with a dimension of $C$, a $7\times 7$ convolution kernel is used for depthwise convolution to obtain the feature map $Fz^{(l-1)}=\text{Concat}\{Fc_{1},Fc_{2},Fc_{3},Fc_{4}\}$. Then, W-MSA or SW-MSA is used to compute the attention map. After the FE-MSA, small objects’ edge and texture features are further enhanced, improving the model’s ability to perceive local and global features. The computation process for each token patch through the RingMo-Aerial block is as follows:
@@
[…]
The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes. Fig.10 presents a comparative analysis of the feature extraction capabilities of the RingMo-Aerial and RingMo models in processing multiview ARS images. The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions. In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features. This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images. These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
@@ 4.3 Pre-training
Figure 10: Comparative analysis of feature heatmaps extracted from multi-view aerial imagery using RingMo-Aerial and RingMo. The feature heatmaps extracted by the RingMo-Aerial model are located in the upper half, while those from the RingMo model are in the lower half of the figure. In these images, the brighter the region, the more the model is focusing on that particular feature, indicating that these areas have higher feature importance for the model.
@@
[…]
@@
[…]
@@
We obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$ through the aforementioned affine transformation process. The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs. By calculating the contrastive loss, we guide the model in learning the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.

### v4

@@ Abstract
Aerial Remote Sensing (ARS) vision tasks present significant challenges due to the unique viewing angle characteristics. Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications. This paper proposes RingMo-Aerial, aiming to fill the gap in foundation model research in the field of ARS vision. A Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism is introduced to strengthen the model’s capacity for small-object representation. Complementarily, an affine transformation-based contrastive learning method improves its adaptability to the tilted viewing angles inherent in ARS tasks. Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and performance in various ARS vision tasks. Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks. This indicates the practicality and efficacy of RingMo-Aerial in enhancing the performance of ARS vision tasks.
@@ 1 Introduction
These challenges highlight the need for specialized intelligent models tailored to ARS applications. While recent foundational models in computer vision and RS have shown potential capabilities, they are often inadequate for ARS due to its unique requirements. These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
@@
In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic ARS foundational models called ARS-Adapter. Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting. Traditional efficient fine-tuning methods, such as Adapter and AdaptFormer, are insufficient in addressing the domain gap between self-supervised pretraining and downstream tasks. This domain gap arises from the inconsistency in feature extraction preferences between pretraining and task-specific objectives. ARS-Adapter mitigates this issue to a certain extent. ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter[6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling. This approach optimizes the performance of the ARS foundational model, making it more adaptable and effective across various ARS vision tasks.
@@
To evaluate the effectiveness of the model, we performed six downstream tasks using 12 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes. The test results show that RingMo-Aerial achieves SOTA performance on all ARS datasets and reaches the leading level on ApRS datasets, which proves its strong generalization ability for intelligent interpretation of aerial and ApRS images. Compared with the single-task model, RingMo-Aerial has apparent advantages in handling multi-tasks in different scenarios.
@@ 2.2 Visual foundation model
Visual foundation models have gained attention as general-purpose solutions for computer vision tasks. Early models, based on convolutional neural networks (CNNs) like VGG [20] and ResNet [21], used convolutional layers for feature extraction and were typically pre-trained on classification tasks using ImageNet [22]. With the advent of Transformers in natural language processing, visual Transformer models like ViT [23], and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training. In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies. At this stage, methods integrating convolution into ViT architectures also began to emerge. Representative works such as CvT [25], Conformer [26], and CoaT [27] enhance local feature modeling by embedding convolution into token projections, using parallel convolution-attention branches, or inserting convolution in early encoder stages. However, most of these approaches rely on stacked 3×3 convolutions, which suffer from a mismatch between the theoretical and effective receptive fields, limiting their ability to capture broader contexts or object-level structures [28]. Besides, several novel architectures, such as ConvNeXt [29] and Mamba[30], have been proposed. However, these newer models have yet to achieve widespread adoption due to technical immaturity.
@@ 2.2.2 Masked Image Modeling
2.2.3 The combination of Contrastive Learning &\& Masked Image Modeling
@@ 3 Problem Description
Figure 5: Similar targets with significant size differences in the same picture. In the Vision Transformer model, objects of different sizes occupy different numbers of tokens. The image is from the VisDrone dataset. The yellow box represents the tokens occupied by the target.
The unique imaging properties of ARS photography stem from the inherent properties of aerial perspective. Specifically, this perspective manifests itself in size differences within the same image due to the tilted angle of the shot. As shown in Fig.5, similar objects at different distances from the ARS are rendered at different sizes in the image. This phenomenon results in distant objects occupying fewer tokens in the Vision Transformer’s input sequence, while closer objects occupy more.
@@
As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\theta$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera.At different horizontal distances, the imaging size corresponding to the same target height differs. Specifically, when the target is directly beneath the camera, its height is not captured in the image. As the horizontal distance increases, the number of pixels occupied by the target’s height initially increases and then decreases. By considering the distances from the target’s lowest and highest points to the camera, given by $x=\sqrt{H^{2}+l^{2}}$ and $y=\sqrt{(H-h)^{2}+l^{2}}$ respectively, we can derive the relationship between the viewing angle $\alpha$ and the horizontal distance $l$ between the target and the camera:
$x^{2}+y^{2}-2xy\cos\theta=h^{2}$ (1)
@@
$\cos\theta=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
By setting $a=H$ and b=Hhb=H-h, we have the target function is:
$f(l)=\cos\theta=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
@@
Since $a=H$ and b=Hhb=H-h, and $a>b$, it follows that ab=h>0a-b=h>0.
@@
This completes the derivation of the maximum value $\theta$ and the corresponding minimum value $\cos\theta$. From Eq. (2), when $l=0$, $\theta$ is equal to 0. From Eqs.(4) and (9), $\theta$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\theta$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\theta$ is again equal to 0. Assuming that the optical axis of the camera is always at the center of the target viewpoint $\theta$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\theta$ of the target. This indicates that if the same target is at different distances from the lens and imaging angles, it will have different representations in the image.
@@ 4.1 Method Overview
The proposed RingMo-Aerial serves as a foundational framework with broad applicability. Its implementation involves three main stages: model design, pre-training, and fine-tuning for downstream tasks. In Section 4.2, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase. To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown in Fig.7. In Section 4.3, we combine the commonly used masked image modeling (MIM) pre-training with additional contrastive learning (CL) methods based on affine transformations during the pre-training stage. This hybrid pre-training approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown in Fig.7. Furthermore, in Section 4.4, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase. The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks.
@@ 4.2 Model Architecture
To address the challenges above, we propose the overall architecture of RingMo-Aerial. As shown in Fig.7 Pre-training Stage 1, considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios. It is an extension to standard Multi-Head Self-Attention (MSA), defined as:
$\mathrm{FE\text{-}MSA}(X)$ $=\mathrm{MSA}(Y_{\mathrm{conv}})$
$=\mathrm{MSA}\bigl{(}X+\mathrm{DWConv}_{7\times 7}(X)\bigr{)}.$ (10)
The input $X$ is first processed through a residual depth-wise $7\times 7$ convolution branch, which expands the local receptive field and compensates for high-frequency components. The enhanced feature map $Y_{\mathrm{conv}}$ is then fed into the MSA layer to enable global information interaction.
@@
The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes. Fig.9 presents a comparative analysis of the feature extraction capabilities of RingMo-Aerial and RingMo models in processing multiview ARS images. The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions. In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features. This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images. These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
In ARS tasks such as object detection and semantic segmentation, small targets are far more prevalent and exhibit more extreme scale variation than in natural or satellite imagery. Standard Vision Transformer MSA or Swin Transformer Window Attention limits each token’s initial receptive field to a single patch, severely hindering edge and detail capture for small targets. Introducing an additional convolutional branch serves to (1) spatially expand the receptive field and (2) bolster the model’s ability to focus attention on fine-scale structures. From a frequency-domain perspective, small targets manifest stronger high-frequency energy, yet attention mechanisms tend to attenuate such components. By inserting a residual $7\times 7$ depth-wise separable convolution prior to attention, FE-MSA enhances band-pass filtering in the high-frequency range, improving detail modeling for small targets.
Fig.7(b) illustrates the RingMo-Aerial Block structure, including two window attention mechanisms optimized by the FE-MSA mechanism. Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block. Taking the first stage as an example, the input feature vector $z^{(l-1)}$ adopts the same partitioning strategy as the standard Swin Transformer. After linear normalization, the feature is passed into the FE-MSA module. In this module, all channels are independently processed using a shared $7\times 7$ depthwise convolution kernel, and the resulting outputs are concatenated along the channel dimension. Then, W-MSA or SW-MSA is used to compute the attention map. After the FE-MSA, small objects’ edge and texture features are further enhanced, improving the model’s ability to perceive local and global features. The computation process for each token patch through the RingMo-Aerial block is as follows:
@@
$\text{FE-MSA}(z^{l-1})=\text{SoftMax}\left(\frac{qk^{T}}{\sqrt{d}}+B\right)v$ (12)
In the $l$-th RingMo-Aerial Block, the input feature map $z^{l-1}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$ is first enhanced via a depthwise convolution in the frequency domain, after which multi-head self-attention is performed within each window: queries $q$, keys $k$, and values $v$ are obtained by linear projections of the enhanced features, their dot-products are scaled by $\sqrt{d}$, augmented with a relative position bias $B$, and normalized by SoftMax, then subsequently weighted-summed with $v$.
@@ 4.3 Pre-training
Figure 9: Comparative analysis of feature heatmaps extracted from multi-view aerial imagery using RingMo-Aerial and RingMo. The feature heatmaps extracted by RingMo-Aerial are located in the upper half, while those from the RingMo model are in the lower half of the figure. In these images, the brighter the region, the more the model is focusing on that particular feature, indicating that these areas have higher feature importance for the model.
@@
$\text{src\_points}=\begin{bmatrix}x_{0}&y_{0}\\ x_{1}&y_{1}\\ x_{2}&y_{2}\\ x_{3}&y_{3}\end{bmatrix},\quad\text{dst\_points}=\begin{bmatrix}u_{0}&v_{0}\\ u_{1}&v_{1}\\ u_{2}&v_{2}\\ u_{3}&v_{3}\end{bmatrix}$ (13)
@@
$\text{dst\_points}=\begin{bmatrix}u_{0}&v_{0}\\ u_{1}&v_{1}\\ u_{2}&v_{2}\\ u_{3}&v_{3}\end{bmatrix}=\begin{bmatrix}x_{0}&y_{0}\\ x_{1}&y_{1}\\ \alpha x_{2}+(1-\alpha)x_{3}&y_{2}\\ \alpha x_{3}+(1-\alpha)x_{2}&y_{3}\end{bmatrix}$ (14)
@@
When generating positive sample pairs via affine transformation, we first apply a random coefficient $\alpha\in[0.05,0.35]$ to compress the bottom width of the image to a specified proportion of its original size while preserving the top width, thereby creating a pixel-level “magnification” effect for distant small objects. The transformed image is then center-cropped to a rectangular region and rescaled to match the model’s input dimensions, effectively removing irrelevant background and further enhancing the visibility of small objects. By sampling $\alpha$ from a diversified range, we achieve multi-scale coverage—from subtle magnification to global micro-distortion—which enables the model to learn semantically consistent embeddings across varying perspectives and spatial resolutions, thereby improving its ability to perceive distant small targets in aerial imagery. For negative samples, we select all full, transformed images from other instances within the same mini-batch.
Specifically, as illustrated in Figure 11, positive samples are generated by first applying an affine transformation that warps the image into an isosceles trapezoid while keeping the top width unchanged. This operation compresses the lower part of the image and stretches the upper part. The transformed image is then cropped to retain only the central rectangular region, which is subsequently resized to form the final positive sample. Although this procedure discards certain areas, it effectively amplifies the upper region of the image—i.e., distant small targets in aerial scenes. Incorporating such positive samples into contrastive pre-training helps the model to better capture scale-invariant features and geometric invariance, thereby improving its ability to detect small objects under complex aerial viewpoints.
We obtain the source image $I_{i}$ and its transformed counterpart $I_{i}^{\prime}$ via the above affine process. The pair $(I_{i},I_{i}^{\prime})$ constitutes a positive sample, while the pairs $(I_{i},I_{j}^{\prime})$ where $i\neq j$, constitute negative samples. By optimizing a contrastive loss, the model is encouraged to learn view- and scale-invariant visual representations, thus enhancing its ability to extract stable physical characteristics from aerial scenes under diverse imaging conditions.
Specifically, the contrastive loss used in our method is defined as follows:
$\mathcal{L}_{\text{contrast}}=-\log\frac{\exp(\text{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{\prime})/\tau)}{\sum\limits_{j=1,j\neq i}^{N}\exp(\text{sim}(\mathbf{h}_{i},\mathbf{h}_{j}^{\prime})/\tau)}$ (18)
where $\mathbf{h}_{i}$ and $\mathbf{h}_{i}^{\prime}$ are the normalized representations of a positive pair, $\text{sim}(\mathbf{h}_{i},\mathbf{h}_{j})=\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\,\|\mathbf{h}_{j}\|}$ is the cosine similarity between two vectors, $\tau$ is a temperature hyper parameter, and NNis the batch size.